# Labelling Data with Unknown References

### v2

@@ Abstract
An evaluator is trustworthy when there exists some agreed-upon way to measure its performance as a labeller. The two ways to establish trustworthiness are either by testing it, or by assuming the evaluator ‘knows’ somehow the way to label the corpus. However, if labelled references (e.g., a development set) are unavailable, neither of these approaches work: the former requires the data, and the latter is an assumption, not evidence. To address this, we introduce an algorithm (the ‘No-Data Algorithm’) by which to establish trust in an evaluator without any existing references. Our algorithm works by successively posing challenges to said evaluator. We show that this is sufficient to establish trustworthiness w.h.p., in such a way that when the evaluator actually knows the way to label the corpus, the No-Data Algorithm accepts its output; and, conversely, flags untrustworthy evaluators when these are unable to prove it. We present formal proofs of correctness and limited experiments.
@@ 1 Introduction
One of the most fundamental problems in machine learning, and AI broadly, is measurement. It has been known for a while that some metrics have a certain level of unreliability; for example, by not corresponding to human judgements appropriately [34, 11, 3, 13]. However, the rise of LLMs as evaluators, typically known as LLMs-as-judges, has led to a substantial divide in the field regarding their trustworthiness as a measurement tool. Their use has also led to broader questions around measurement in AI; specifically, about the validity of some results [11, 33, 25].
@@
A key component in the ‘measurement in AI question’, and the object of study of our work, is whether the tool providing the evaluation (the evaluator) may be trusted. This has become more critical given the fast adoption of foundation models in industry and scientific research, and the continuous need to scale. In these cases, trust is usually established based on some sort of ground truth: an existing set of labelled references by which to perform statistical analyses (e.g., a $t$-test). However, it is often the case that said references do not necessarily exist: perhaps the problem is too niche, or the annotation too costly. It follows then that the only way one can establish trust on an unknown evaluator in a no-reference scenario is via formal methods (i.e., formal proofs).
@@
In this paper we introduce an algorithm (which we call the ‘No-Data Algorithm’) by which to establish trust in an evaluator without previously-labelled data. The algorithm is designed to work in-line with contemporary machine learning techniques, where an evaluator claims that it knows the map between datapoints and labels. It is up to a verifier to decide, based on the outcome of certain challenges, whether the evaluator may be trusted. The No-Data Algorithm requires a linear number of calls and does not require labelled datasets (e.g., development or training sets) or knowledge of the labels in the test set to establish trust.
We show that our algorithm, after $r$ calls, bounds confidence on the labeller to $(1/4)^{r}$. This success rate reflects the ‘true’ accuracy of the evaluator: if the evaluator knows how to label the set, the accuracy and success rates will be high; if the evaluator lies about knowing the labels, the accuracy could be high, but the success rate won’t. We present small experiments in synthetic data for traditional machine-learning (i.e., a decision tree) and contemporary (r. an LLM) methods which adjust to the theory’s predictions, although we note our contribution is theoretical in nature. [^1] Code and data may be found in https://github.com/adewynter/no_data_algorithm. Amongst our experimental findings we note that the behaviour of a reliable (read: trained) evaluator adjusts to our theoretical predictions, while an LLM requires more work due to prompting and spurious reasoning. Nonetheless, when properly tuned, it adjusts to the predictions. Our results demonstrate that trust in an evaluator may be established via formal methods without the need for references. A side effect of our experimentation opens a wide range of questions around faithful generation over arbitrary strings for a generative model.
@@ 2 Related Work
There has been considerable work on attempting to create reliable evaluators. Most of contemporary work relies on LLMs and various prompting strategies or call stacks, although the problem of (un)reliability/trustworthiness of an evaluator is older than generative models. For example, it has been known that the use of BLEU for tasks other than English-based machine translation is not accurate [22, 28]; and even then, that it does not correspond well to human judgements [34, 11].
@@
More generally, it is known that metric choice, and hence the choice of evaluator, is system-dependent [4, 25, 10, 39, 32]. Poorly-developed rubrics also lead to unreliable human judgements [7, 20], and thus the evaluator trustworthiness problem is also a rubric problem. There are algorithms [27, 47] and paradigms [26] for learning with noisy teachers. Likewise, provable methods have been used with success in other areas, such as representation learning [15], and to improve robustness via synthetic data generation [8, 9]. However, they all assume that a (perhaps noisily) labelled dataset exists.
@@
Relevant to our work, the use of LLMs as evaluators has promised extraordinary scaling capabilities, such as fast data generation and evaluation without humans [29, 6, 5, 23, 41, 48]. Some of these criticisms are that they do not correlate with human judgements well [3, 13, 4, 45, 37]; are sensitive to their prompt [24, 14, 46]; their performance could be the product of memorisation [19, 44, 35]; and even that they do not understand the task at all [43, 40, 46]. Their reasoning capabilities have been put into question, from the results being strongly dependent on the choice of metrics [36, 4], to findings that their output reasoning steps contain spurious reasoning [38, 18, 42]. Most of the work on improving trust, however, relies on prompting strategies or call stacks to other LLMs [21].
@@
The main subroutine from the No-Data Algorithm, which we call the Evaluator-Verifier (EV) protocol is based off the well-known Arthur-Merlin (AM) protocol from Babai [2]. It is a type of zero-knowledge proof; see [12] and [1] for primers on the subject. That said, the EV protocol is different in many regards to the AM protocol, since it is designed to fit arbitrary inputs and the larger No-Data Algorithm. In particular, the EV protocol always reveals its secrets, which sets it apart from applications in theoretical computer science. Likewise, our use of the rubric and the aggregator functions are generalisations to partial functions that are not present (or needed) in previous works.
3 Background
4 Notation
Our work relates to datapoints $x\in X$ and labels $y\in Y$. Whenever the label set $Y$ is unspecified, it is the binary label set $Y=\{0,1\}$. We also let $X\subset\{0,1\}^{n}$; that is, every datapoint is a $n$-bit binary string modelling a phenomenon. The assumptions on $X$ and $Y$ are to simplify our proofs; see Section 9 for a discussion on their relevancy.
Likewise, for a binary string $x\in\{0,1\}^{n}$, we denote its relevant subset set as $S_{x}\subset 2^{x}$. The construction of this specific set will depend on the setup, and will be discussed in detail in Section 6.1.
For two functions $f\colon B\rightarrow C,g\colon A\rightarrow B$ we write the composition $f\circ g\colon A\rightarrow C$ as $fg$.
For two sets $A$, $B$, we denote equality up to isomorphism as $A\cong B$, where $A\cong B$ if $f(A)=f(B)$ but $A\neq B$ for some $f\colon U\rightarrow V$. We also consider the special case of equality up to permutation over subsets as $A\equiv B$, where $A\equiv B$ if $\forall a\in S_{A},\exists b\in S_{B}\,.\,g(a)=g(b)$, for some fixed function $g\colon U\rightarrow V$. We assume all functions to be deterministic throughout, unless stated otherwise.
@@ 5 Definitions and Problem Setup
In line with standard learning theory, we assume that there exists an unknown map $f\colon X\rightarrow Y$ that takes datapoints to labels faithfully–that is, $f$ provides the true labelling. Our task is to determine whether a function $E\colon X\rightarrow Y$ (a labeller, or evaluator) is equivalent to $f$, without any knowledge of the map itself.
For that, we introduce four definitions:
A criterion $c\colon X\rightarrow\{0,1\}$ is a map that takes in a datapoint $X$ and maps it to a single bit.
The rubric $C\colon\times_{i,\dots,n}\{0,1\}\rightarrow\{0,1\}^{n}$ takes in the binary outputs from an ordered set of criteria $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ and returns a bitstring that is the concatenation of these outputs: $C=c_{1}|c_{2}|\dots|c_{n}$. We say a rubric is total if it explicitly decomposes nonlinear criteria with arity $>1$ (e.g. the xor $c_{i}=c_{a}\oplus c_{b}$), and denote it as $\bar{\mathcal{C}}$. When the evaluation of said criteria can be done separately (e.g., by testing first $c_{a}$, then $c_{b}$, and then their xor $c_{a}\oplus c_{b}$), we say it is a total evaluation.
The aggregator $\sigma\colon\{0,1\}^{n}\rightarrow Y$ is a function that takes in a binary string of length $n$ (the output from the rubric) and maps it to the label set.
@@
Intuitively, the rubric is the set of reasons why a datapoint $x$ would take a given label (for example, an annotation rubric passed in to human labellers, the last layer in a model, etc.); while the aggregator decides the final label based on that rubric (e.g., via an activation function, a majority vote, etc.).
@@
The map $f\colon X\rightarrow Y$ is a composition of the rubric and the aggregator. That is,
@@
While not a particular strong assumption (after all, datapoints must have a reason to acquire a label and, when there is more than one reason, there must be a way to decide how to aggregate them) this decomposition is crucial to our work.
@@ 6 The No-Data Algorithm
The No-Data Algorithm is an algorithm with two components (players): the evaluator and the verifier. [^2] In theoretical computer science, the evaluator is typically known as the prover. We use this more-intuitive naming to highlight the distinction between our work (data-labelling) and others (e.g. theorem-proving). Throughout the algorithm, for every datapoint $x\in X$, the evaluator must convince the verifier that it is able to label $x$ correctly (i.e., that it knows $f$).
For this, the players run a multi-round sub-game which we refer to as the Evaluator-Verifier (EV) protocol. The EV protocol takes in $x$, and either succeeds or ends in failure, w.h.p., based on challenges posed by the verifier on a $x^{\prime}$ similar to $x$ generated by the evaluator. It then returns the proposed label and the status (success or failure).
The algorithm then, in case of failure, decides to flip the label with probability $\phi$. At the end of the loop, it aggregates the results of the EV protocol and returns the final labelling and success count. The final labelling depends on the choice of $\phi$, and the evaluator’s trustworthiness (and hence that of said labelling) is be determined by measuring the success count.
The full algorithm may be found in Algorithm 1. Given that the EV protocol is a fundamental part of the No-Data Algorithm, we focus the remainder of this section to describe it, and provide a characterisation of the No-Data Algorithm (bounds, correctness, etc.) in Section 7.1.
@@
15:  return  $predictions$, $successes$
Algorithm 1 The No-Data Algorithm. It works by calling a subroutine (the EV Protocol) to provably determine the trustworthiness of the labels produced by the evaluator. Based on its output (success or failure), the algorithm then determines which label to assign to the corpus, with a stochastic label flip $\phi$ added to it in case of failures.
@@ 6.1 The Evaluator-Verifier (EV) Protocol
In the EV protocol both players are given a datapoint $x\in X$, and the rubric $C$, but nothing else. In particular, $\sigma$ is not explicitly given to either player. However, the evaluator claims to know the map $f\colon X\rightarrow Y$. Since $C$ is given, then the claim is that it has prior knowledge of $\sigma$. The implication–were the evaluator to convince the verifier–then is that its choice of label $y$ for $x$ may be trusted.
Remark that the verifier does not know if $y$ is correct. In order to be convinced, the verifier chooses one out of two challenges, neither of which rely on $\sigma$ or $y$. Based on whether the evaluator passes or fails them, the verifier builds its own belief on the evaluator’s trustworthiness.
The entire point of the EV protocol (and thus of the No-Data Algorithm) is for the evaluator to convince the verifier that its choice of label for $x$ is correct. If the evaluator fails, it follows that the correct label for $x$ could be the opposite label. Hence, regardless of success, at the end of the EV protocol the evaluator reveals its label. See Figure 1 for a diagram of the EV protocol. We now describe each of the components and flow.
@@ 6.1.1 Setup
At every round of the EV protocol, the evaluator takes in $x$ and generates a similar datapoint $x^{\prime}$ and a ‘partial’ label $\tilde{y}^{\prime}$ that–ideally–is equal to $y$. It is partial because there is no access to $f$, and hence the best the evaluator can do is to generate a guess $\tilde{y}^{\prime}$ based on the encoding $C(x^{\prime})$. If the evaluator indeed knows $f$, an application of $\sigma$ to $C(x^{\prime})$ would lead to the ‘true’ label $y$.
That said, once it has generated the tuple $\langle x^{\prime},\tilde{y}^{\prime}\rangle$, the evaluator may not change them in this round.
Remark that isomorphism is not a sufficient condition for $\tilde{y}^{\prime}=y$ since $C$ might not be total. Hence, both structural and valuation checks are needed. Therefore, after the tuple has been generated, the verifier then chooses uniformly at random one of the two following challenges:
Figure 1: Diagram of the EV protocol. At every round, the evaluator (blue) generates a datapoint $x^{\prime}$ similar to $x$, and a partial label $\tilde{y}^{\prime}$. It then responds to one of two (chosen uniformly at random) challenges by the verifier (orange). The game is repeated by $r$ rounds and returns a succeeded state, unless there is a failure in the challenges, and thus returns a failed state. In either case, it also returns the proposed partial label, $\tilde{y}^{\prime}$.
@@
Setup: The verifier tests the evaluator by asking it to show a similar point that leads to the same total evaluation of $C$. This challenge is surrogated by testing whether $x\equiv x^{\prime}$ over $\bar{\mathcal{C}}$.
@@
namely, whether there exists at least one substring in $x^{\prime}$ matching the (total) evaluation of a criterion from $x$.
@@
Setup: The verifier asks for the rationale as to why $f(x^{\prime})=y$. Since $y$ is not available, it checks whether $x^{\prime}\cong x$.
@@
6.1.2 Verification
If the evaluator failed the tests, the EV Protocol returns a failure state. Otherwise, it restarts from Section 6.1.1 over and over until the maximum number of rounds $r$ has been reached, or the evaluator fails a test. In either case, at the end of the call the EV Protocol also returns $\tilde{y}^{\prime}$.
@@ 6.1.2 Verification
Since the challenges are public, the evaluator could cheat and generate an $x^{\prime}$ to pass one of the challenges, but not both. Otherwise, the implication is that it knows $f$, and it will pass the challenges every time. This is why the challenges assert both structure ($x\equiv x^{\prime}$) and final encoding ($x\cong x^{\prime}$). This observation will be elaborated further in our correctness proofs ( Section 7.1).
Remark 2.
The check for Challenge 1 (demanding equality up to criterion) is very strict. We discuss this in Section 9 and present a relaxed extension in Appendix C.
7 Algorithm Characterisation
In this section we provide proofs of correctness (Section 7.1). Runtime bounds may be found in Appendix B.
7.1 Correctness

### v3

@@ Abstract
An evaluator is trustworthy when there exists some agreed-upon way to measure its performance as a labeller. The two ways to establish trustworthiness are either by testing it, or by assuming the evaluator ‘knows’ somehow the way to label the corpus. However, if labelled references (e.g., a development set) are unavailable, neither of these approaches work: the former requires the data, and the latter is an assumption, not evidence. To address this, we introduce an algorithm (the ‘No-Data Algorithm’) by which to establish trust in an evaluator without any existing references. Our algorithm works by successively posing challenges to said evaluator. We show that this is sufficient to establish trustworthiness w.h.p., in such a way that when the evaluator actually knows the way to label the corpus, the No-Data Algorithm accepts its output; and, conversely, flags untrustworthy evaluators when these are unable to prove it. We present formal proofs of correctness, empirical tests, and applications to LLMs-as-judges on low-resource languages.
@@ 1 Introduction
One of the most fundamental problems in AI is the question of measurement. It has been known for a while that some metrics have a certain level of unreliability; for example, by not corresponding to human judgements appropriately (Reiter, 2018; Gehrmann et al., 2023; Bavaresco et al., 2024; Hada et al., 2024). With the rise of LLMs as evaluators, typically known as LLMs-as-judges, scaling measurement to multiple subproblems has become ubiquitous (de Wynter, 2025). However, their reliability and trustworthiness as a measurement tool has led to a substantial divide in the research community, and specifically around the validity of some results (Gehrmann et al., 2023; Rabinovich and Anaby Tavor, 2025; Moghe et al., 2023).
@@
A key component of the question of measurement, and the focus of our work, is whether the tool providing the evaluation (the evaluator) may be trusted. This trust is typically established with some sort of ground truth, such as an existing set of labelled references with which to perform statistical analyses (e.g., a $t$-test), or a benchmark with previously-published results. Due to uniqueness of the problem, cost of annotation, or other reasons, references may not readily exist, or be too scarce to be statistically significant. Likewise, contamination of public benchmarks has become a major concern (see, e.g., Sainz et al. 2023). It follows then that the only way one can scientifically establish trust on an evaluator whose capabilities are unknown, in a no-reference scenario, is with a formal proof.
@@
In this paper we introduce an algorithm (the ‘No-Data Algorithm’) by which to establish trust in an evaluator without previously-labelled data. The algorithm is designed to work in-line with contemporary machine learning techniques–namely, LLMs-as-judges–where an evaluator claims it knows the map between datapoints and labels. It is up to a verifier to decide, based on the outcome of certain challenges, whether the evaluator may be trusted. The No-Data Algorithm requires a linear number of calls and does not require labelled datasets (e.g., development or training sets) or knowledge of the labels in the test set to establish trust.
We show that our algorithm, after $r$ calls, bounds confidence on the evaluator to a $(1/4)^{r}$ probability of error. This success rate reflects the evaluator’s knowledge–or lack thereof: if it knows how to label the set, the accuracy and success rate, as reported by the No-Data Algorithm, will be high. Conversely, if the evaluator lies about knowing the labels, the accuracy could be high–perhaps by luck–but the success rate will not. We present empirical tests with traditional machine-learning (i.e., a decision tree) and contemporary (r. LLMs) evaluation methods. We also showcase an application of the No-Data Algorithm by judging an LLM-as-a-judge’s ability to label a dataset in West Frisian, a low-resource language. Both experiments adjust to the theory’s predictions. [^1] Code and data may be found in https://github.com/adewynter/no_data_algorithm. Our results demonstrate that trust in an evaluator may be established via formal methods without the need for references, and enables their application in scenarios where labels may be extremely scarce, such as market research, low-resource languages (the subject of our experiments), and the medical domain.
@@ 2 Related Work
There has been considerable work on attempting to create reliable evaluators. Most of contemporary work relies on LLMs and various prompting strategies or call stacks, although the problem of (un)reliability/trustworthiness of an evaluator is older than generative models. For example, it has been known that the use of BLEU for tasks other than English-based machine translation is not accurate (Liu et al., 2016; Novikova et al., 2017); and even then, that it does not correspond well to human judgements (Reiter, 2018; Gehrmann et al., 2023).
@@
More generally, it is known that metric choice, and hence the choice of evaluator, is system-dependent (Chen et al., 2024; Moghe et al., 2023; Flamich et al., 2025; von Däniken et al., 2024; Pombal et al., 2025). Poorly-developed rubrics also lead to unreliable human judgements (Clark et al., 2021; van der Lee et al., 2019), and thus the evaluator trustworthiness problem is also a rubric problem. There are algorithms (Northcutt et al., 2017; Zhang and Sabuncu, 2018) and paradigms (Northcutt et al., 2021) for learning with noisy teachers. Likewise, provable methods have been used with success in other areas, such as representation learning (Jovanović et al., 2023), and to improve robustness via synthetic data generation (Dimitrov et al., 2022; Fischer et al., 2022). However, they all assume that a (perhaps noisily) labelled dataset exists.
@@
Relevant to our work, the use of LLMs as evaluators has promised extraordinary scaling capabilities, such as fast data generation and evaluation without humans (Open AI, 2023; Chiang et al., 2024; Chiang and Lee, 2023; Liu et al., 2023; Wei et al., 2024; Zheng et al., 2023). Some of these criticisms are that they do not correlate with human judgements well (Bavaresco et al., 2024; Hada et al., 2024; Chen et al., 2024; de Wynter et al., 2025; Tjuatja et al., 2024); are sensitive to their prompt (Lu et al., 2022; Hida et al., 2024; Ye and Durrett, 2022); their performance could be the product of memorisation (Lee et al., 2023; de Wynter et al., 2023; Sainz et al., 2023); and even that they do not understand the task at all (de Wynter and Yuan, 2024; Webson and Pavlick, 2022; Ye and Durrett, 2022). Their reasoning capabilities have been put into question, from the results being strongly dependent on the choice of metrics (Schaeffer et al., 2023; Chen et al., 2024), to findings that their output reasoning steps contain spurious reasoning (Turpin et al., 2023; Lanham et al., 2023; Wu et al., 2024). Most of the work on improving trust, however, relies on prompting strategies or call stacks to other LLMs (Li et al., 2024).
@@
The main subroutine from the No-Data Algorithm, which we call the Evaluator-Verifier (EV) protocol is based off the well-known Arthur-Merlin (AM) protocol from Babai (1985). It is a type of zero-knowledge proof; see Goldreich (2009) and Arora and Barak (2009) for primers on the subject. That said, the EV protocol is different in many regards to the AM protocol, since it is designed to fit arbitrary inputs and the larger No-Data Algorithm. In particular, the EV protocol always reveals its secrets, which sets it apart from applications in theoretical computer science. Likewise, our use of the rubric and the aggregator functions are generalisations to partial functions that are not present (or needed) in previous works.
@@ 3 Definitions and Problem Setup
3.1 Notation
Our work relates to datapoints $x\in X$ and labels $y\in Y$. Whenever the label set $Y$ is unspecified, it is the binary label set $Y=\{0,1\}$. We also let $X\subset\{0,1\}^{n}$; that is, every datapoint is a $n$-bit binary string modelling a phenomenon. The assumptions on $X$ and $Y$ are to simplify our proofs; see Section 7 for a discussion on their relevancy. For a binary string $x\in\{0,1\}^{n}$, we denote its relevant subset set as $S_{x}\subset 2^{x}$. The construction of this specific set will depend on the setup, and will be discussed in detail in Section 4.1.
For two functions $f\colon B\rightarrow C,g\colon A\rightarrow B$ we write the composition $f\circ g\colon A\rightarrow C$ as $fg$. For two sets $A$, $B$, we denote equality up to isomorphism as $A\cong B$, where $A\cong B$ if $f(A)=f(B)$ but $A\neq B$ for some $f\colon U\rightarrow V$. We also consider the special case of equality up to permutation over subsets as $A\equiv B$, where $A\equiv B$ if $\forall a\in S_{A},\exists b\in S_{B}\,.\,g(a)=g(b)$, for some fixed function $g\colon U\rightarrow V$. We assume all functions to be deterministic throughout, unless stated otherwise.
3.2 Definitions
In line with standard learning theory (Valiant, 1984), we assume that there exists an unknown map $f\colon X\rightarrow Y$ that takes datapoints to labels faithfully–that is, $f$ provides the true labelling. Our task is to determine whether a function $E\colon X\rightarrow Y$ (a labeller, or evaluator) is equivalent to $f$, without any knowledge of the map itself. This is measured with its error rate, or, conversely, its accuracy.
We also introduce four definitions which we will use in our work:
A criterion $c\colon X\rightarrow\{0,1\}$ maps a datapoint $X$ to a single bit.
A rubric $C\colon\times_{i,\dots,n}\{0,1\}\rightarrow\{0,1\}^{n}$ maps the binary outputs from an ordered set of criteria $\mathcal{C}=\{c_{1},\dots,c_{n}\}$ to a bitstring that is the concatenation of these outputs: $C=c_{1}|c_{2}|\dots|c_{n}$. We say a rubric is total if it explicitly decomposes nonlinear criteria with arity >1>1 (e.g. the xor $c_{i}=c_{a}\oplus c_{b}$), and write it as $\bar{\mathcal{C}}$. When the evaluation of said criteria can be done separately (e.g., by testing first $c_{a}$, then $c_{b}$, and then their xor $c_{a}\oplus c_{b}$), we say it is a total evaluation.
The aggregator $\sigma\colon\{0,1\}^{n}\rightarrow Y$ is a function that maps $n$-bit binary strings (rubric outputs) to $Y$.
@@ 3.2 Definitions
Intuitively, the rubric is the set of reasons why a datapoint $x$ would take a given label (for example, an annotation rubric passed in to human labellers, the last layer in a model, etc.); while the aggregator decides the final label based on that rubric (e.g., via an activation function, a majority vote, etc.). Natural-language examples are in Appendix E.
@@
The map $f\colon X\rightarrow Y$ is a composition of the rubric and the aggregator:
@@
It is not a strong assumption: after all, datapoints must have a reason to acquire a label. When there is more than one reason, there must be a way to decide how to aggregate them. However, this decomposition is crucial to our work.
@@ 4 The No-Data Algorithm
The No-Data Algorithm is an algorithm with two components (players): the evaluator and the verifier.Throughout the algorithm, for every datapoint $x\in X$, the evaluator must convince the verifier that it is able to label $x$ correctly (i.e., that it knows $f$).
For this, the players run a multi-round sub-game which we refer to as the Evaluator-Verifier (EV) protocol. The EV protocol takes in $x$, and either succeeds or fails, w.h.p., based on challenges posed by the verifier on a $x^{\prime}$ similar to $x$ generated by the evaluator. The EV protocol returns the proposed label and the status (success or failure).
The No-Data Algorithm then flips the label with probability $\phi$ in case of failure. After observing every $x\in X$, it returns the success count and final labelling. The evaluator’s trustworthiness (and hence the final labelling’s) is established with the success count.
The full algorithm is in Algorithm 1. We and provide a characterisation of the No-Data Algorithm (bounds, correctness, etc.) in Section 5.
@@
15:return  predictionspredictions, successessuccesses
Algorithm 1 The No-Data Algorithm. For every $x\in X$ it calls the EV protocol to determine the trustworthiness of the label returned by the evaluator, $y=E(x)$. Based on its output, the algorithm flips $y$ w.p. $\phi$ in case of failure. Finally, it returns the success count and the labels.
@@ 4.1 The Evaluator-Verifier (EV) Protocol
In the EV protocol both players are given a datapoint $x\in X$, rubric $C$, and nothing else. In particular, $\sigma$ is not explicitly given to either player. The evaluator claims to know the map $f\colon X\rightarrow Y$. Since $C$ is given, the claim is then that it has prior knowledge of $\sigma$. The verifier’s goal is to be sufficiently sure that the evaluator’s choice of label $x$ for a given $x\in X$ may be trusted.
@@ 4.1.1 Setup
The EV protocol is played for $r$ rounds. At every round, the evaluator generates a similar datapoint $x^{\prime}$ to $x$, and a ‘partial’ label $\tilde{y}^{\prime}$ that–supposedly–is equal to $y$. It is partial because there is no access to $f$, and hence the best the evaluator can do is to guess $\tilde{y}^{\prime}$ based on $C(x^{\prime})$. Once it has generated the tuple $\langle x^{\prime},\tilde{y}^{\prime}\rangle$, the evaluator may not change it in this round. The verifier poses one of two challenges, selected uniformly at random. Either challenge measures equality between $x^{\prime}$ and $x$, either up to isomorphism or up to permutation. If the evaluator fails, the protocol terminates and returns failure. Otherwise, it starts again. Either in success or failure, the protocol also returns $\tilde{y}^{\prime}$. See Figure 1 for a diagram of the flow of the EV protocol.
The verifier challenges are central to the work, and perform both structural and valuation checks:
Figure 1: EV protocol flow. At every round, the evaluator (blue) generates an $x^{\prime}$ similar to $x$, and a partial label $\tilde{y}^{\prime}$. It then answers one of two (chosen uniformly at random) challenges by the verifier (orange). If the evaluator does not pass the challenge, the protocol returns failure. Otherwise, the game is repeated. If the rounds are over, it returns a succeeded state. In either case, it also returns $\tilde{y}^{\prime}$.
@@
Setup: The verifier tests the evaluator by asking it to show a similar point that leads to the same total evaluation of $C$, namely, whether $x\equiv x^{\prime}$ over $\bar{\mathcal{C}}$.
@@
that is, if there is at least one substring in $x^{\prime}$ matching the (total) evaluation of a criterion from $x$.
@@
Setup: The verifier asks why $f(x^{\prime})=y$. Since $y$ is not available, it checks whether $x^{\prime}\cong x$.
@@
Since the challenges are public, the evaluator could cheat and generate an $x^{\prime}$ to pass one of the challenges, but not both. Otherwise, the implication is that it knows $f$, and it will pass the challenges every time. This is why the challenges assert both structure ($x\equiv x^{\prime}$) and final encoding ($x\cong x^{\prime}$). This observation is central to our correctness proofs (Section 5).
5 Characterisation